A 75.0 kW, 230-V shunt generator has a generated emf of 243.5 V. If the field current is 12.5 A at
rated output, what is the armature resistance? (Ans. 0.0399).
1
Expert's answer
2020-12-18T11:11:11-0500

Answer

Shunt current

"I_s=\frac{P}{V}=\frac{75000}{230}=326A"


So "R_s=\frac{V}{I}=\frac{230}{326}=0.7\Omega"

So using formula

"I_s=\frac{V}{R_s+R_A}"

Where

RA is armature resistance. Putting all value then we get

RA=0.0399"\Omega"